OpenStudy (anonymous):

Find the Laplace transform of the function: f(t) =Sin 2t if π ≤ t ≤ 2π; f(t) = 0 if t <π or if t>2π

OpenStudy (anonymous):

Rewrite using the unit step functions.

OpenStudy (anonymous):

So, we have: \[ F(s)=\int_{0}^\infty e^{-st}\sin(t)\;dt \]For \(t\in [\pi, 2\pi]\) And, we have: \[ F(s)=0 \]For \(t\in \mathbb{R}\setminus [\pi, 2\pi]\) Evaluate this case to receive the answer.

OpenStudy (anonymous):

I'm still having a hard time. I just started learning this material last Friday.

OpenStudy (anonymous):

What problems are you having?

OpenStudy (anonymous):

how to even go about starting the problem